What is the molecular formula of each compound? (Type your answer using the format C6H12O6 for C6H12O6 and use the same order of the elements as in the empirical formulas.) (a) empirical formula C3H6O2 ( = 74.08 g/mol) (b) empirical formula HgCl ( = 472.1 g/mol) (c) empirical formula NO2 ( = 92.02 g/mol) (d) empirical formula CHCl ( = 96.94 g/mol)
Answer #1

a) empirical formula C3H6O2 ( = 74.08 g/mol)

empirical formula mass = 3 x 12 + 6 x 1 + 2 x 16 = 74

n = 74.08 / 74 = 1

molecular formula = n x empirical formula

                              = 1 x C3H6O2

                             = C3H6O2

molecular formula = C3H6O2

(b) empirical formula HgCl ( = 472.1 g/mol)

HgCl mass = 200.59 + 35.5 = 236.09

n = 472.1 / 236.09 = 2

molecular formula = 2 x HgCl

                               = Hg2Cl2

molecular formula = Hg2Cl2

    (c) empirical formula NO2 ( = 92.02 g/mol)

molecular formula =N2O4

(d) empirical formula CHCl ( = 96.94 g/mol)

CHCl = 12 + 1 + 35.5 = 48.5

n = 2

molecular formula = C2H2Cl2